# Accelerated Rotation-Invariant Convolution for UAV Image Segmentation

### v1

@@ V-D Post-Segmentation Processing and Visualization
The rotation invariant U‑Net segmentation network was trained to classify plant species in the imagery, and species‑label outputs from all frames (see Fig. 12) were georeferenced and mosaicked into a seamless orthomosaic. The resulting species‑labeled mosaic was exported to Google Earth Pro for visualization (Fig. 13) and for analysis of species distribution across the field.
@@
Figure 13: Geo-referenced P1 RGB imagery mosaic (a) and corresponding Geo-referenced mask mosaic (b) displayed in Google Earth Pro.
As shown in Fig. 14, the plant area consists of 51.41% clover leaf, 37.89% grass, 3.13% dock leaf, and 7.56% weeds and other vegetation, indicating clover as the dominant ground cover followed by grass, with minimal presence of dock leaves and miscellaneous weeds.

### v2

@@ V-D Post-Segmentation Processing and Visualization
The rotation-invariant U-Net segmentation network was trained to segment plant species in the imagery, and the plant-species segmentation masks from all frames (see Fig. 12) were subsequently georeferenced and mosaicked into a seamless orthomosaic. The resulting species‑labeled mosaic was exported to Google Earth Pro for visualization (Fig. 13) and for analysis of species distribution across the field. As shown in Fig. 14, the plant area consists of 51.41% clover leaf, 37.89% grass, 3.13% dock leaf, and 7.56% weeds and other vegetation, indicating clover as the dominant ground cover followed by grass, with minimal presence of dock leaves and miscellaneous weeds.
@@
Figure 13: Geo-referenced P1 RGB imagery mosaic (Top) and corresponding Geo-referenced mask mosaic (Bottom) displayed in Google Earth Pro.